A sample of 347 subscribers to Wired magazine shows the mean time spent using the Internet is 12 hours per week, with a sample standard deviation of 6.1 hours.

Find the 80% confidence interval for the mean time Wired subscribers spend on the Internet. (Round the final answers to 2 decimal places.)

The confidence interval is between  and  .
Answer #1

Solution :

Given that,

= 12

s =6.1

n = Degrees of freedom = df = n - 1 =347 - 1 =346

a ) At 80% confidence level the t is ,

= 1 - 80% = 1 - 0.80 = 0.20

  / 2= 0.20 / 2 = 0.10

t /2,df = t0.10,346 =: 1.284 ( using student t table)

Margin of error = E = t/2,df * (s /n)

= : 1.284* (6.1 / 346)

= 0.42

The 80% confidence interval is,

- E < < + E

12 -0.42 < <12 +0.42

11.58 < < 12.42